# MS-0001-1922.65

### v1

@@ 4.3 Lagrangian Relaxation
    The optimal solution of the Lagrangian Primal Problem $(\ell^{*}_{L-P},h^{*}_{L-P})$ can be computed as:
    $\ell^{*}_{L-P}\;=$ $\;\inf_{\ell\;\in\;[\ell(\xi^{as}),\;h(\xi^{ru})]}\left\{\lambda\;\mathbb{E}[\mathbbm{1}(\delta^{+}\leq\ell)]-\gamma_{1}\;\phi(\ell)+\gamma_{2}\;\ell\right\},$
    $h^{*}_{L-P}\;=$ $\;\inf_{h\;\in\;[\ell(\xi^{as}),\;h(\xi^{ru})]}\left\{(1-\lambda)\;\mathbb{E}[\mathbbm{1}(\delta^{-}\geq h)]+\gamma_{1}\;\phi(h)-\gamma_{2}\;h\right\}.$
@@
    Let the optimal value of the objective function of the Primal Problem and the Lagrangian Primal Problem be $\mathcal{Z}^{*}_{P}$ and $\mathcal{Z}^{*}_{L-P}(\gamma_{1},\gamma_{2})$, respectively. Then, we have:
    $\mathcal{Z}^{*}_{L-P}(\gamma_{1},\gamma_{2})\;\leq\;\mathcal{Z}^{*}_{P}.$
Proposition 4.4 states that the Lagrangian Primal Problem provides a lower bound for the optimal value function for given Lagrange multipliers. Now it remains to identify Lagrange multipliers that provide the tightest possible lower bound. This can be done by solving the following dual problem:
Lagrangian Dual Problem
$\max_{\gamma_{1},\gamma_{2}}\;$ $\min_{\ell,h}\;\;\mathcal{L}(\ell,h,\gamma_{1},\gamma_{2})$
s.t. $\;\;h\;\leq\;h(\xi^{ru})$
$\;\;\ell\;\geq\;\ell(\xi^{as})$
[…]
Let $\mathcal{Z}^{*}_{L-D}$ be the optimal objective value function of the Lagrangian Dual Problem. By weak duality, we have, $\mathcal{Z}^{*}_{L-D}\;\leq\;\mathcal{Z}^{*}_{P}$. Using this fact, we propose a simple and efficient algorithm (Algorithm 1) to iteratively solve the dual problem and find a tight lower bound. We next describe this algorithm.
Algorithm 1 Lagrangian Relaxation Algorithm
[…]
2:Initialize lower bound $\mathcal{Z}^{(1)}_{LB}=10^{-3}$ and upper bound $\mathcal{Z}^{(1)}_{UB}=\mathcal{Z}_{P}(\ell^{(1)},h^{(1)})$.
3:Initialize Lagrange multipliers $\gamma^{(1)}_{1}=0$ and $\gamma^{(1)}_{2}=0$.
[…]
5:     Compute $(\ell^{*}_{L-P},h^{*}_{L-P})$ by solving the following two sub-problems:
$\ell^{*}_{L-P}\;=$ $\;\inf_{\ell\;\in\;[\ell(\xi^{as}),\;h(\xi^{ru})]}\left\{\lambda\;\mathbb{E}[\mathbbm{1}(\delta^{+}\leq\ell)]-\gamma^{(n)}_{1}\;\phi(\ell)+\gamma^{(n)}_{2}\;\ell\right\},$
$h^{*}_{L-P}\;=$ $\;\inf_{h\;\in\;[\ell(\xi^{as}),\;h(\xi^{ru})]}\left\{(1-\lambda)\;\mathbb{E}[\mathbbm{1}(\delta^{-}\geq h)]+\gamma_{1}^{(n)}\;\phi(h)-\gamma_{2}^{(n)}\;h\right\}.$
6:     Set $\mathcal{Z}_{L-P}(\gamma^{(n)}_{1},\gamma^{(n)}_{2})=\lambda\;\mathbb{E}[\mathbbm{1}(\delta^{+}\leq\ell^{*}_{L-P})]+(1-\lambda)\;\mathbb{E}[\mathbbm{1}(\delta^{-}\geq h^{*}_{L-P})]+\gamma^{(n)}_{1}\,g^{(n)}_{1}+\gamma^{(n)}_{2}\,g^{(n)}_{2},$     where $g^{(n)}_{1}=\phi(h^{*}_{L-P})-\phi(\ell^{*}_{L-P})-p$ and $g^{(n)}_{2}=(\ell^{*}_{L-P}-h^{*}_{L-P})$.
7:     Set $\mathcal{Z}^{(n+1)}_{LB}=\max\left\{\mathcal{Z}^{(n)}_{LB},\;\mathcal{Z}_{L-P}(\gamma^{(n)}_{1},\gamma^{(n)}_{2})\right\}$.
8:     if $g_{1}^{(n)}\leq 0$ and $g_{2}^{(n)}\leq 0$ then
9:      Set $\mathcal{Z}^{(n+1)}_{UB}=\min\left\{\mathcal{Z}^{(n)}_{UB},\;\lambda\;\mathbb{E}[\mathbbm{1}(\delta^{+}\leq\ell^{*}_{L-P})]+(1-\lambda)\;\mathbb{E}[\mathbbm{1}(\delta^{-}\geq h^{*}_{L-P})]\right\}$.
10:      if $\lambda\;\mathbb{E}[\mathbbm{1}(\delta^{+}\leq\ell^{*}_{L-P})]+(1-\lambda)\;\mathbb{E}[\mathbbm{1}(\delta^{-}\geq h^{*}_{L-P})]\;\leq\;\mathcal{Z}^{(n)}_{UB}$ then
11:       Set $\ell^{(\text{best})}=\ell^{*}_{L-P}$ and $h^{(\text{best})}=h^{*}_{L-P}$.
12:     else
13:      Set $\mathcal{Z}^{(n+1)}_{UB}=\mathcal{Z}^{(n)}_{UB}$.
14:     if $(\mathcal{Z}^{(n+1)}_{UB}-\mathcal{Z}^{(n+1)}_{LB})/\mathcal{Z}^{(n+1)}_{LB}\;\leq\;\epsilon$ then
15:      Stop and return the best solution $(\ell^{(\text{best})},\;h^{(\text{best})})$.
16:     Update Lagrange multipliers:
$\gamma^{(n+1)}_{1}=\max\left\{0,\gamma^{(n)}_{1}+\alpha^{(n)}\;g^{(n)}_{1}\right\},\quad\gamma^{(n+1)}_{2}=\max\left\{0,\gamma^{(n)}_{2}+\alpha^{(n)}\;g^{(n)}_{2}\right\},$
  where $\alpha^{(n)}=c\;(\mathcal{Z}^{(1)}_{UB}-\mathcal{Z}_{L-P}(\gamma^{(n)}_{1},\gamma^{(n)}_{2}))/\left\lVert(g^{(n)}_{1},g^{(n)}_{2})\right\rVert^{2}$ and $c\in(0,2]$.
[…]
Utilizing the structural properties outlined in §4.2 and exploited in Algorithm 1, we present our main theoretical result, which introduces a systematic approach to solving the optimization problem (1)-(5).
Theorem 4.5
Let $\phi(y)=\mathbb{P}(f(X)\leq y)$ be the empirical CDF of $f(X)$. Then, for the optimization problem (1)-(5), we have:
  1. (a)
    If $h(\xi^{ru})\;>\;\ell(\xi^{as})$ and $\phi(h(\xi^{ru}))-\phi(\ell(\xi^{as}))\;\leq\;p$, then $(\ell(\xi^{as}),h(\xi^{ru}))$ is the unique optimal solution.
  2. (b)
    If $h(\xi^{ru})\;>\;\ell(\xi^{as})$ and $\phi(h(\xi^{ru}))-\phi(\ell(\xi^{as}))\;>\;p$, then an approximate solution can be obtained by Algorithm 1.
@@
    If $h(\xi^{ru})\;<\;\ell(\xi^{as})$, then the problem is infeasible.
@@ 5 Policy Recommendation and Impact Evaluation
Using our results from the previous section, we now propose a data-driven clearance policy that leverages a human-algorithm approach to assist the FDA in its decision-making. We also leverage the dataset we have assembled (see Section 2) and investigate the effectiveness of our policy vis-a-vis the FDA’s current practice. To this end, we randomly split the data into training data (70%) and testing data (30%), and we use the ML model presented in 3.2 to provide the inputs required for the optimization procedures discussed in the previous section (see Algorithm 1).
When making use of our policy, three parameters should be set by the decision-maker: $\xi^{ru}$, $\xi^{as}$, and $p$, which correspond to the rates of rejection of unsafe devices, the acceptance of safe devices, and the upper bound on the workload, respectively. We employ a cross-validation approach to investigate the impact of these input parameters on various metrics. This can assist decision-makers in selecting the right input parameters that align with their criteria. Table 5 summarizes our results by showing the impact of these parameters on the acceptance and rejection rates of both safe and unsafe devices as well as the FDA’s workload which currently stands at 100% as all devices are evaluated by the FDA’s committees. In this table, each input parameter is considered at three levels: low (L), medium (M), and high (H). For the parameters $\xi^{ru}$ and $\xi^{as}$, we have $L=0.3$, $M=0.5$, and $H=0.7$. For the parameter $p$, the values are $L=0.4$, $M=0.6$, and $H=0.8$.
To provide a clear performance evaluation, we consider a conservative scenario where we assume that the risk labels assigned to deferred devices do not contribute to enhancing the evaluation process conducted by FDA committees. Specifically, we assume that our policy achieves the same recall rate as the current practice for all deferred devices. Thus, our reported estimates of potential impact are conservative since it is likely that the risk labels assigned can themselves enable the FDA’s human experts to improve their decisions.
As an example, consider the case where $\xi^{ru}=M$, $\xi^{as}=L$, and $p=M$ (i.e., M-L-M combination). We observe that using our policy results in rejecting 50.1% of unsafe devices and accepting 30.7% of safe devices. Among the accepted devices, 8.3% would experience a future recall, while 19.9% of the rejected devices would face no future recall. It is worth noting that the rejection rate of unsafe devices, reported in Table 5, can also be interpreted as the percentage improvement in the recall rate achieved by the proposed policy in comparison to the recall rate under the FDA’s current practice. This interpretation is valid because the number of unsafe devices rejected under our policy is equivalent to the difference between the number of unsafe devices accepted under the FDA’s current practice and the number of unsafe devices accepted under our policy. Accordingly, this policy leads to 50.1% improvement in the recall rate compared to the FDA’s current practice. Additionally, it results in a 51.5% reduction in the workload of the FDA’s committees, because only 48.5% of the devices will be forwarded to the FDA’s committees for in-depth evaluation and the remaining 51.5% will be automatically accepted or rejected.
From Table 5, we observe that there is an interplay between the threshold parameters and their influence on the acceptance and rejection rates for safe and unsafe devices. A higher value of $\xi^{ru}$ leads to a more stringent policy, resulting in an increased rate of rejection for safe devices. Similarly, a higher value of $\xi^{as}$ is associated with a corresponding increase in the acceptance rate of unsafe devices. This highlights the intricate balance required between safety and efficiency in the policy. Additionally, the optimality gap shown in the last column of Table 5 is very low, showcasing our Lagrangian-based algorithm’s ability in discovering near-optimal solutions. Finally, we note that the “N/A” values in the last three rows indicate that the corresponding input parameters render the problem infeasible. The infeasibility of these cases stems from the fulfillment of condition (c) in Theorem 4.3.
We observe substantial gains in certain scenarios, such as the L-L-L combination, which results in a remarkable 74.1% workload reduction and a significant 72.2% recall rate improvement (rejection of unsafe devices). However, it is essential to note a vital caveat. Some combinations that offer significant improvements in workload reduction and recall rate also lead to an unreasonably high rate of rejection of safe devices. This phenomenon arises in the specific setting of low and high thresholds that, in certain cases, reject a substantial proportion of safe devices. Therefore, the selection of input parameters must strike a delicate balance between reducing workload and maintaining an acceptable rate of rejection for safe devices.
Table 5: Comparative analysis of key metrics across combinations of input parameters
{tblr}
hline1-2,29 = -, $\xi^{ru}$ $\xi^{as}$ $p$ Reject
Unsafe % Accept
Safe % Accept
Unsafe % Reject
Safe % Workload % Gap %
L L L 72.2 30.7 8.3 42.6 25.9 23.7
L L M 34.8 30.7 8.3 9.9 59.1 2.0
L L H 29.9 30.7 8.3 7.6 61.7 0.0
L M L 35.7 50.4 22.5 10.2 39.6 0.4
L M M 29.9 50.4 22.5 7.6 42.6 0.0
L M H 29.9 50.4 22.5 7.6 42.6 0.0
L H L 29.9 69.3 38.0 7.6 24.1 0.0
L H M 29.9 69.3 38.0 7.6 24.1 0.0
L H H 29.9 69.3 38.0 7.6 24.1 0.0
M L L 72.6 30.7 8.3 44.4 24.3 26.2
M L M 50.1 30.7 8.3 19.9 48.5 0.0
M L H 50.1 30.7 8.3 19.9 48.5 0.0
M M L 50.1 50.4 22.5 19.9 29.4 0.0
M M M 50.1 50.4 22.5 19.9 29.4 0.0
M M H 50.1 50.4 22.5 19.9 29.4 0.0
M H L 50.1 69.3 38.0 19.9 10.9 0.0
M H M 50.1 69.3 38.0 19.9 10.9 0.0
M H H 50.1 69.3 38.0 19.9 10.9 0.0
H L L 69.5 30.7 8.3 38.9 29.5 0.0
H L M 69.5 30.7 8.3 38.9 29.5 0.0
H L H 69.5 30.7 8.3 38.9 29.5 0.0
H M L 69.5 50.4 22.5 38.9 10.3 0.0
H M M 69.5 50.4 22.5 38.9 10.3 0.0
H M H 69.5 50.4 22.5 38.9 10.3 0.0
H H L “N/A” “N/A” “N/A” “N/A” “N/A” “N/A”
H H M “N/A” “N/A” “N/A” “N/A” “N/A” “N/A”
H H H “N/A” “N/A” “N/A” “N/A” “N/A” “N/A”
@@
A Representative Policy. To gain deeper insights, we next focus on the L-L-M combination as a representative setting, and investigate various aspects of the resulted policy. Figure 7 illustrates the optimized low and high thresholds, as well as the distribution of predicated risk values for unrecalled and recalled devices. The right-hatched region to the left of the low threshold represents the proportion of safe devices accepted by the policy. Similarly, the left-hatched region to the right side of the high threshold represents the proportion of unsafe devices rejected by the policy. Additionally, the overlapping hatched region to the left of the low threshold depicts the proportion of devices falsely accepted, while the overlapping hatched region to the right of the high threshold depicts the proportion of devices falsely rejected.
@@
Table 6 shows the impact of the representative policy. As it can be seen, using this policy leads to a 43.0% reduction in the workload of the FDA’s committees and a 38.9% improvement in the recall rate percentage. Investigating the policy, we observe that applicant devices with a recall risk estimated to be below 0.16 are accepted, while applicant devices with a recall risk estimated to be higher than 0.46 are rejected. The other applicant devices are categorized by a risk label and deferred for judgment by an FDA’s committee of human experts. Under our conservative evaluation, following this policy results in the rejection of 38.9% of unsafe devices and the acceptance of 28.8% of safe devices. Among the accepted devices, 7.9% will experience a recall in the future, while 13.7% of the rejected devices will have no recalls in the future.
@@
0.16 0.46 43.0 38.9
A sensitivity analysis is also conducted to assess the impact of the assumption that risk labels do not enhance the evaluation process conducted by FDA committees. The results highlight the further improvements that risk labels can offer (see Section 9 for details).
@@ 5.1 Policy Impact: Costs
In the previous sections, we evaluated the impact of our proposed policy in terms of measures such as correct acceptance and rejection rates as well as the resulted workload. In this section, we examine its impact on costs. To this end, we note that a number of independent studies have shown that recalled or prematurely failed devices have likely cost Medicare billions of dollars in recall-related health care expenditures (HHS 2017). In one year alone, the FDA received reports of nearly three-thousand potential device-related deaths, over one-hundred thousand potential device-related injuries, and over two-hundred thousand adverse event reports concerning medical devices (Zuckerman et al. 2011). Nonetheless, the full extent of injury to patients that is attributable to recalled devices is unknown. Thus, we conservatively assess the cost-savings conforming to the replacement costs of a recalled device. However, it is worth noting that the true cost of a recall may far exceed our calculations (i.e., the replacement costs incurred by the manufacturers) because it will also include potential cost of related injuries. In 2016, for example, a settlement of various state and federal personal injury litigations from recalled pelvic mesh products totaled \$121 million (Medtronic 2023). In a separate case in 2016, personal injury claims concerning a bone graft product were settled for \$26 million (Medtronic 2016). The fact that these amounts reflect settlements, not affirmative decisions by courts, suggests that full compensation for personal injury could have been even higher.
To our knowledge, there are no previous scholarly publications that have systematically assessed the costs of recalling the spectrum of medical devices. Our review of the literature uncovered two notable publications on this topic. The first publication was a 2017 report by the Office of Inspector General (OIG). Consistent with our own literature review, this report noted that there is no reliable up-to-date estimate of Medicare costs associated with recalled medical devices (HHS 2017). The second publication was a white paper published by the McKinsey Center for Government and cited by the FDA in a presentation discussing the benefit of reducing medical device failure cost (Tack 2021). McKinsey estimated that non-routine events “such as major observations, recalls, warning letters, and consent decrees, along with associated warranties and lawsuits” cost the industry between \$2 billion and \$5 billion per year on average. The total cost includes \$1.5 billion to \$3 billion per year on non-routine costs, plus \$1 billion to \$2 billion in lost sales of new and existing products. This suggests that annual non-routine costs of recalls can range between \$0.5 billion to \$3 billion per year (Fuhr et al. 2013).
@@
Our assessment of the replacement costs for recalled devices is based on administrative claims data titled “Medicare Durable Medical Equipment, Devices & Supplies” (MDMEDS) for the years 2013-2020. This data is published annually by the Centers for Medicare and Medicaid Services (CMS) and contains information on usage, payments, and submitted charges organized by National Provider Identifier (NPI) and Healthcare Common Procedure Coding System (HCPCS) code (CMS 2021b). The dataset is based on information gathered from CMS administrative claims data for Original Medicare Part B beneficiaries available from the CMS Chronic Conditions Data Warehouse. The data are summarized from 100% final-action Durable Medical Equipment, Prosthetic, Orthotics and Supplies (DMEPOS) non-institutional claim line items (CMS 2021b).
We calculate the replacement costs by medical specialty using the HCPCS codes and descriptions reported in the MDMEDS data. As noted by the OIG report, the replacement cost of recalled medical devices cannot be tracked to individual devices solely with claims data as Medicare claim forms do not contain a field for reporting medical device-specific information (HHS 2017). Consequently, our estimate of replacement costs is specific to the medical specialty, not individual devices. We determined the medical specialty for over 99% of the 1,351 unique HCPCS codes/descriptions in the MDMEDS 2013-2020 data by first identifying keywords in the device classification names available in the 510(k) submission data for each medical specialty and then matching those keywords to the HCPCS descriptions in the MDMEDS data. For example, an HCPCS description containing the word “ostomy,” which is a procedure used to treat various diseases of the urinary or digestive systems, was classified as having the medical specialty Gastroenterology/Urology. In a slightly more intricate case, HCPCS descriptions containing the words “glucose monitor” or “glucose” and “monitor” were classified as having the medical specialty Clinical Chemistry. A detailed table showing the crosswalk between keywords and medical specialties is available in Table 10.
Once the medical specialties for the HCPCS codes were established, we calculated the average Medicare allowed amount per medical specialty. This calculation was based on the total supplier claims, which reflects the number of products ordered by the referring provider. The Medicare allowed amount includes the amount Medicare paid, the deductible and coinsurance amounts owed by the beneficiary, as well as any amount owed by a third-party payer (CMS 2021a).
We were able to calculate the average Medicare allowed amount for over 75% of the devices in our test dataset based on their respective medical specialties. However, for the devices for which we were unable to compute the average Medicare allowed amount, we faced a challenge in assigning a specific medical specialty to them. Among them, Radiology devices account for 13.7% of all devices in the test dataset and the remaining specialties cumulatively account for less than 8.5% of devices. We believe that the chief reason we could not calculate Medicare costs for these specialties is that the underlying 510(k) devices are not single-use devices, expended on a single patient. In the case of radiology, we examined the majority of 510(k) cleared products between 2013 and 2020 and determined that these devices were either imaging software or multi-use equipment used in providing radiology services such as radiosurgery or radiotherapy. For the purposes of our impact assessment, we created a low and high estimate of the average Medicare allowed amount when we could not directly calculate the average Medicare allowed amount. The low estimate is the lowest average Medicare allowed amount across all of the specialties. The high estimate is the weighted average Medicare allowed amount across all specialties, with weights derived from the frequency of each medical specialty in our testing dataset.
@@
Finally, we calculated the potential cost-savings from implementing our data-driven policy derived from the L-L-M combination (see Figure 7) on the testing set with 9,582 medical devices. In Step 1, we identified 317 recalled devices that were cleared by the FDA but our policy would have rejected if implemented. In Step 2, we determined the number of units being recalled for each of the aforementioned devices using the FDA recall data, which amounted to approximately 64.8 million units. In Step 3, we used the low and high average Medicare allowed amounts to determine the low and high estimated cost-savings that would have occurred had these devices been rejected pursuant to our proposed policy.
@@
Figure 8 shows the percentage of recalls avoided by our proposed policy per medical specialty in the testing set (Step 1). For example, 55.1% for the GU medical specialty indicates that 109 recalls out of 198 recalls corresponding to the GU medical specialty in the testing set were avoided by our proposed policy. As can be seen, Radiology (RA) devices have the highest frequency, followed by Anesthesiology (AN) and Hematology (HE). On the other spectrum, Obstetrics/Gynecology (OB), Clinical Toxicology (TX), and Physical Medicine (PM) have the lowest frequency (zero). Comparing with the recall rate per medical specialty in our dataset (2008-2020), we observe that OB, PM, and TX are among the top four medical specialties in terms of having low recall rates. The alignment between low avoided recalls and low overall recall rates suggests that our policy is functioning as intended, particularly in areas where it is most needed. Table 7 reports the total average cost-savings per medical specialty (Step 3). Among them, Orthopedic (OR), Cardiovascular (CV), and Gastroenterology & Urology (GU) constitute the top three medical specialties with the highest average cost savings, respectively. The results based on our testing set show that the overall cost-savings from implementing our policy range between \$7.7 billion and \$8.7 billion during the several years in our study period. Given that there are approximately 3,000 510(k) submissions annually (Dubin et al. 2021 and Kadakia et al. 2023), by extrapolating from the total cost-savings in our testing set, a rough estimation of the resulted annual cost-savings is between \$2.4 billion and \$2.7 billion.
@@
Table 7: Total cost-savings associated with six most cost-saving medical specialties in the testing set
@@
cell24 = fg=OxfordBlue, cell34 = fg=OxfordBlue, cell44 = fg=OxfordBlue, cell74 = fg=OxfordBlue, hline1-2,8 = -, Medical specialty & Recalls Avoided % Average
Product Quantity Average Cost-Savings

### v2

@@ 4.3 Nested Search
    If $h(\xi^{ru})\;>\;\ell(\xi^{as})$ and the workload constraint (14) holds for $\ell\,=\,\ell(\xi^{as})$ and $h\,=\,h(\xi^{ru})$, then $(\ell(\xi^{as}),h(\xi^{ru}))$ is the unique optimal solution,
@@
    If $h(\xi^{ru})\;>\;\ell(\xi^{as})$ and the workload constraint (14) does not hold for $\ell\,=\,\ell(\xi^{as})$ and $h\,=\,h(\xi^{ru})$, then for any given $\epsilon>0$, our nested search algorithm returns an $\epsilon$-optimal solution after at most
    $\mathcal{O}\left(\frac{\bar{L}\left(h(\xi^{ru})-\ell(\xi^{as})\right)}{\epsilon}\;\log_{2}\left(\frac{2\bar{L}_{\phi}(1-\lambda)\left(h(\xi^{ru})-\ell(\xi^{as})\right)}{\epsilon}\right)\right)$
    iterations, where $\lambda\in(0,1)$, $\bar{L}_{\phi}$ and $\bar{L}_{h}$ are, respectively, upper bounds on the common Lipschitz constant of ϕ±()superscriptitalic-$\phi^{\pm}(\cdot)$ and the Lipschitz constant of $h^{*}(\ell)$, and $\bar{L}=\bar{L}_{\phi}\left(\lambda+(1-\lambda)\bar{L}_{h}\right)$. Setting $\Delta\;=\;\epsilon/\bar{L}$ and $\zeta\;=\;\epsilon/2(1-\lambda)\bar{L}_{\phi}$ guarantees $\epsilon$-optimality and ensures that the workload constraint is satisfied within a tolerance of $\epsilon/2(1-\lambda)$,
@@
    If $h(\xi^{ru})\;<\;\ell(\xi^{as})$, then the problem is infeasible,
@@ 5 Policy Recommendation and Impact Evaluation
Using our results from the previous section, we now propose a data-driven clearance policy that leverages a human-algorithm approach to assist the FDA in its decision-making. We also leverage the dataset we have assembled (see Section 2) and investigate the effectiveness of our policy vis-a-vis the FDA’s current practice. To this end, we randomly split the data into training data (70%) and testing data (30%), and we use the ML model presented in 3.1 to provide the inputs required for the optimization procedures discussed in the previous section.
When making use of our policy, three parameters should be set by the decision-maker: $\xi^{ru}$, $\xi^{as}$, and $p$, which correspond to the rates of rejection of unsafe devices, the acceptance of safe devices, and the upper bound on the workload, respectively. This can assist decision-makers in selecting the right input parameters that align with their criteria. We investigate the impact of these parameters on the acceptance and rejection rates of both safe and unsafe devices as well as the FDA’s workload, which currently stands at 100% as all devices are evaluated by the FDA’s committees. To provide a clear performance evaluation, we consider a conservative scenario where we assume that the additional information for deferred devices do not contribute to enhancing the evaluation process conducted by FDA committees. For any device that our policy defers to the FDA committee, we assume the committee takes the same action it took historically, and we evaluate safety using that device’s observed recall outcome. Thus, our reported estimates of potential impact are conservative since it is likely that the additional information for deferred devices can themselves enable the FDA’s human experts to improve their decisions. To do so, each input parameter is considered at three levels: low (L), medium (M), and high (H). For the parameters $\xi^{ru}$ and $\xi^{as}$, we have $L=0.3$, $M=0.5$, and $H=0.7$. For the parameter $p$, the values are $L=0.4$, $M=0.6$, and $H=0.8$. A summary of results is presented in Appendix 13. The results indicate that there is an interplay between the threshold parameters and their influence on the acceptance and rejection rates for safe and unsafe devices. A higher value of $\xi^{ru}$ leads to a more stringent policy, resulting in an increased rate of rejection for safe devices. Similarly, a higher value of $\xi^{as}$ is associated with a corresponding increase in the acceptance rate of unsafe devices. This highlights the intricate balance required between safety and efficiency in the policy.
@@
To gain deeper insights, we next focus on the L-L-M combination as a representative setting, and investigate various aspects of the resulting policy. Figure 9 illustrates the optimized low and high thresholds, as well as the distribution of predicted risk values for unrecalled and recalled devices. The right-hatched region (\) to the left of the low threshold represents the proportion of safe devices correctly accepted by the policy, while the right-hatched region (\) to the right of the high threshold represents the proportion of safe devices falsely rejected. Similarly, the left-hatched region (/) to the right side of the high threshold represents the proportion of unsafe devices correctly rejected by the policy, while the left-hatched region (/) to the left of the low threshold represents the proportion of unsafe devices falsely accepted. Additionally, the overlapping hatched region between the low and high thresholds shows the proportion of deferred devices.
@@
Table 4 shows the impact of the representative policy in the testing set. As can be seen, using this policy leads to a 40.5% reduction in the workload of the FDA’s committees and a 32.9% improvement in the recall rate percentage. Investigating the policy, we observe that applicant devices with a recall risk estimated to be below 0.060 are accepted, while applicant devices with a recall risk estimated to be higher than 0.177 are rejected. The other applicant devices are deferred for judgment by an FDA committee of human experts. Under our conservative evaluation, following this policy results in the rejection of 32.9% of unsafe devices and the acceptance of 30.7% of safe devices. Among the accepted devices, 8.9% will experience a recall in the future, while 9.7% of the rejected devices will have no recalls in the future.
@@
0.060 0.177 40.5 32.9
A robustness check is also conducted to assess the potential impact of selection bias that may arise from the absence of information on devices rejected by the FDA. Our findings suggest that including such devices does not undermine the overall effectiveness of our proposed policy (see Appendix 11 for details).
@@ 5.1 Policy Impact: Costs
In the previous sections, we evaluated the impact of our proposed policy in terms of measures such as correct acceptance and rejection rates as well as the resulting workload. In this section, we examine its impact on costs. To this end, we note that a number of independent studies have shown that recalled or prematurely failed devices have likely cost Medicare billions of dollars in recall-related health care expenditures (HHS 2017). In one year alone, the FDA received reports of nearly three-thousand potential device-related deaths, over one-hundred thousand potential device-related injuries, and over two-hundred thousand adverse event reports concerning medical devices (Zuckerman et al. 2011). Nonetheless, the full extent of injury to patients that is attributable to recalled devices is unknown. Thus, we conservatively assess the cost savings conforming to the replacement costs of a recalled device. However, it is worth noting that the true cost of a recall may far exceed our calculations (i.e., the replacement costs incurred by the manufacturers) because it will also include the potential cost of related injuries. In 2016, for example, a settlement of various state and federal personal injury litigations from recalled pelvic mesh products totaled \$121 million (Medtronic 2023). In a separate case in 2016, personal injury claims concerning a bone graft product were settled for \$26 million (Medtronic 2016). The fact that these amounts reflect settlements, not affirmative decisions by courts, suggests that full compensation for personal injury could have been even higher. McKinsey estimated that non-routine events “such as major observations, recalls, warning letters, and consent decrees, along with associated warranties and lawsuits” cost the industry between \$2 billion and \$5 billion per year on average. The total cost includes \$1.5 billion to \$3 billion per year on non-routine costs, plus \$1 billion to \$2 billion in lost sales of new and existing products. This suggests that annual non-routine costs of recalls can range between \$0.5 billion to \$3 billion per year (Fuhr et al. 2013).
The focus of our cost savings analysis is on payer savings, as measured by the avoidance of replacement costs associated with device recalls. While our policy reduces recall-related expenditures, it also results in some false rejections related to cases in which a device with no observed recall in our dataset is recommended for rejection. Although false rejections can incur costs for manufacturers (e.g., additional delays, resubmissions), the cost of such false rejections for 510(k) devices is relatively small for patients or payers. Because 510(k) devices must be substantially equivalent to an already-marketed device, most do not provide one-of-a-kind treatments. Thus, if our policy erroneously rejects certain 510(k) devices, near-identical substitutes (i.e., the predicate or a similar cleared device) presumably remain on the market. Therefore, the additional harm to patients or payers is minimal clinically and financially.
@@
Our assessment of the replacement costs for recalled devices is based on administrative claims data titled “Medicare Durable Medical Equipment, Devices & Supplies” (MDMEDS) for the years 2013-2020. This data is published annually by the Centers for Medicare and Medicaid Services (CMS) and contains information on usage, payments, and submitted charges organized by National Provider Identifier (NPI) and Healthcare Common Procedure Coding System (HCPCS) code (CMS 2021). The dataset is based on information gathered from CMS administrative claims data for Original Medicare Part B beneficiaries available from the CMS Chronic Conditions Data Warehouse. The data are summarized from 100% final-action Durable Medical Equipment, Prosthetic, Orthotics and Supplies (DMEPOS) non-institutional claim line items (CMS 2021). We calculate the replacement costs by medical specialty using the HCPCS codes and descriptions reported in the MDMEDS data. The details of this calculation, which requires creating a crosswalk between keywords and medical specialties, are discussed in Appendix 14.
@@
We calculated the potential cost savings from implementing our data-driven policy derived from the L-L-M combination (see Figure 9) on the testing set with 9,572 medical devices. In Step 1, we identified 325 recalled devices that were cleared by the FDA but our policy would have rejected if implemented. In Step 2, we determined the number of units being recalled for each of the aforementioned devices using the FDA recall data, which amounted to approximately 64.8 million units. In Step 3, we used the low and high average Medicare allowed amounts to determine the low and high estimated cost savings that would have occurred had these devices been rejected pursuant to our proposed policy.
@@
Figure 10 shows the percentage of recalls avoided by our proposed policy per medical specialty in the testing set (Step 1). For example, 29.1% for the CV medical specialty indicates that 44 recalls out of 151 recalls corresponding to the CV medical specialty in the testing set were avoided by our proposed policy. As can be seen, Immunology (IM) devices have the highest frequency, followed by Clinical Chemistry (CH) and Radiology (RA). On the other spectrum, Dental (DE), Ear, Nose, & Throat (EN), Obstetrics/Gynecology (OB), and Physical Medicine (PM) have the lowest frequency (zero). Comparing with the recall rate per medical specialty in our dataset, we observe that EN, OB, and PM are among the top four medical specialties with low recall rates (less than 1%) and DE has a recall rate of 1.9%. The alignment between low avoided recalls and low overall recall rates suggests that our policy is functioning as intended, particularly in areas where it is most needed. Table 5 reports the total average cost savings associated with four most cost-saving medical specialties (Step 3). Among them, Orthopedic (OR) is the medical specialty with the highest average cost savings. The results based on our testing set show that the overall cost savings from implementing our policy range between \$5.5 billion and \$5.6 billion during the several years in our study period. Given that there are approximately 3,000 510(k) submissions annually (Dubin et al. 2021 and Kadakia et al. 2023), by extrapolating from the total cost savings in our testing set, a rough estimation of the resulted annual cost savings is \$1.7 billion.
@@
Table 5: Total cost savings associated with four most cost-saving medical specialties in the testing set
@@
hline1-2,6 = -, Medical specialty & Recalls Avoided % Average
Product Quantity Average cost savings